OpenStudy (anonymous):

Given the equation Square root of 2x plus 1 = 3, solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

@ganeshie8 help plz?

ganeshie8 (ganeshie8):

\(\large \sqrt{2x+1} = 3\)

ganeshie8 (ganeshie8):

Is the equation like above ?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

I got 4 as the answer idk if its right

ganeshie8 (ganeshie8):

then weesy :) the moment u see a radical, squaring should strike in ur mind

OpenStudy (anonymous):

also idk if its extraneous or not

ganeshie8 (ganeshie8):

4 is right !

OpenStudy (anonymous):

yes ! Okay and it is extraneous ? I believe

ganeshie8 (ganeshie8):

to knw if its extraneous or not, do below :- 1) put x = 4 in original given equation 2) if it satisfies, then its NOT an extraneous solution

OpenStudy (anonymous):

so its not

OpenStudy (anonymous):

in this case

OpenStudy (anonymous):

\[\sqrt[-4]{x-3}=12\] solve for x and identify if it is an extraneous solution. thats another one I need help with

ganeshie8 (ganeshie8):

not sure how to approach this :|

OpenStudy (anonymous):

so is it 4 extraneous or not?